OpenStudy (anonymous):

What is the sum of the first 11 terms of the arithmetic series in which a1 = 12 and d = 5? (Points : 1) 38 549 143 209

OpenStudy (anonymous):

im thinking its 209

OpenStudy (anonymous):

Wolfram-Alpha provides the formula \[na _{1} + nd(n-1)/2\] 11(12)+11(5)(10)/2 = 407 which is 'none of the above'. I get the same result from a little Python script. So I'm suspicious of those numbers. But I have been wrong many times before now.

OpenStudy (anonymous):

What is the sum of the first 11 terms of the arithmetic series in which a1 = 12 and d = 5? first we need to find the nth term of this series for this use the formula an= a1 + (n-1)d so an =12 +(11-1)5 an= 62 and to find the sum use the formula = [n(a1+an) ]/2 = [11(12+62)]/2 = 11*74/ 2 = 11 *37 = 407